# Smallest Suffixient Sets: Effectiveness, Resilience, and Calculation

### v4

@@ 4.3 Reversals
    aba and aab.
Similarly, let $w_{3}^{R}=\text{\tt\textdollar}_{2}\texttt{baa}\text{\tt\#}_{2}\texttt{baa}\texttt{c}\text{\tt\textdollar}_{1}\texttt{aba}\text{\tt\#}_{1}\texttt{aba}\texttt{c}\text{\tt\textdollar}_{0}\texttt{aab}\text{\tt\#}_{0}\texttt{aab}\texttt{c}$. The super-maximal right-extensions of $w_{3}^{R}$ are:
@@
    baa and $\text{\tt\textdollar}_{2}$;
@@
    $\texttt{baa}\text{\tt\#}_{2}$ and baac; $\texttt{aba}\text{\tt\#}_{1}$ and abac; $\texttt{aab}\text{\tt\#}_{0}$ and aabc;
@@
    $\texttt{ac}\text{\tt\textdollar}_{0}$; $\texttt{ac}\text{\tt\textdollar}_{1}$;
@@
    aba and aab.
One can see that $\mathtt{sre}(w_{3})=14$, $\mathtt{sre}(w_{3}^{R})=12$, and hence, $\mathtt{sre}(w_{3})-\mathtt{sre}(w_{3}^{R})=2$, as stated in Lemma 7.
4.4 Lower and upper bounds on the sensitivity of $\chi$
Formally, the additive sensitivity of a measure of repetitiveness $\mu$ to a string operation $\rho$ can be defined as a function $AS_{\mu,\rho}:\mathbb{Z}^{+}\rightarrow\mathbb{R}$, where $AS_{\mu,\rho}(n)=\max_{w\in\Sigma^{n}}(\max_{w^{\prime}\in\rho(w)}(\mu(w^{\prime}))-\mu(w))$, that is, the maximum achievable difference among all the strings. Overall, we obtain the following result on the additive sensitivity of $\mathtt{sre}$, which, by Corollary 1, can be written in terms of $\chi$.
Corollary 2
The following bounds on the additive sensitivity of the measure $\chi$ to string operations hold:
@@ 4.4 Lower and upper bounds on the sensitivity of $\chi$
    $AS_{\chi,\rho}(n)=\Omega(\log n)$ for $\rho\in\{\mathtt{ins},\mathtt{del},\mathtt{sub},\mathcal{R}(\cdot)\}$;
@@
    $AS_{\chi,\mathtt{rev}}(n)=\Omega(\sqrt{n})$, where $\mathtt{rev}(w)=\{w^{R}\}$.
Proof
Claim 1 follows by Lemma 6, where $n=|w|=\Theta(2^{k})$ and $AS_{\chi,\rho}(n)=\Omega(k)=\Omega(\log n)$, for all $\rho\in\{\mathtt{ins},\mathtt{del},\mathtt{sub},\mathcal{R}(\cdot)\}$. Claim 2 follows by Lemma 7, where $n=|w_{k}|=\Theta(k^{2})$ and $AS_{\chi,\mathtt{rev}}(n)=\Omega(k)=\Omega(\sqrt{n})$. ∎
Lemma 8
Proof
5 On the Multiplicative Sensitivity of $\mathtt{sre}$ to Reverse
We prove in this section that $\chi$ at most doubles when we reverse the string.
In the following, we denote the set of symbols preceding a substring $x$ (or left-extensions of $x$) of $w$ by $L_{x}(w)$, and we use $l_{x}(w)$ to denote an arbitrary element of $L_{x}(w)$ (we drop $w$ if it is clear from the context). For simplicity of analysis, we assume that $w$ contains a special symbol \$ at the beginning and at the end. This does not matter in asymptotic terms, as $\mathtt{sre}(w)=\mathtt{sre}(\text{\tt\textdollar}w\text{\tt\textdollar})\pm O(1)$, and similarly, $\mathtt{sre}(\text{\tt\textdollar}w^{R}\text{\tt\textdollar})=\mathtt{sre}(w^{R})\pm O(1)$. We do this to (i) ensure all right-maximal substrings of $w$ have at least two left-extensions, and hence, right-extensions have at least one left-extension, and (ii) to provide the necessary symmetry to prove Lemma 10 below.
@@ 5 On the Multiplicative Sensitivity of $\mathtt{sre}$ to Reverse
Please note that left-extensions are not analogous to right-extensions. A right-extension is a string $xa$ such that $xb$ also occurs for some $b\neq a$, while a left-extension of $x$ is just any symbol preceding $x$. A left-extension can be unique, whereas right-extensions come (at least) in pairs. We will rely on the following observation.
Observation 5.1
If $xa$ is a super-maximal right-extension, then $L_{xa}\cap L_{xb}=\emptyset$ for any $b\neq a$.
Proof
@@
Otherwise, $xa$ is not super-maximal, because there exists $c\in L_{xa}\cap L_{xb}$, for which $cxa\in\mathcal{F}_{w}$ and $cxb\in\mathcal{F}_{w}$, and thus $xa$ is a proper suffix of the right-extension cxacxa of the right-maximal string $cx$. ∎
@@
Let us introduce a special tree of super-maximal extensions, which will be helpful to prove our result.
Definition 7
@@
The prefix super-maximal right-extension tree (PSR tree) of $w$ is a compacted trie of all the super-maximal right-extensions of $w$. Since some of those strings can be prefixes of others, internal nodes of the trie may also represent strings, and their corresponding nodes are not compacted. Distinct leaves do not represent prefixes of each other.
The following result is key for our proof.
Lemma 9
@@
Let $w$ be a string starting with a unique symbol \$. If $x$ is a right-maximal substring of $w$, and $xa$ is a super-maximal right-extension of $x$ in $w$, then $x^{R}$ is a right-maximal substring of $w^{R}$.
Proof
@@
Because $xa$ is a right-extension in $w$, there exists $xb\in\mathcal{F}_{w}$ with $b\neq a$. Any symbol preceding $xa$ cannot precede $xb$, otherwise, $xa$ would not be super-maximal. Since $w$ starts with the unique symbol \$, if $x\neq\varepsilon$, then $xa$ and $xb$ have occurrences that are not prefixes of $w$, and hence, they are preceded by different symbols. It follows $x^{R}$ is right-maximal in $w^{R}$.∎
Lemma 10
@@
    If $x$ is a proper prefix of $y$, then and $y$ has to start with $xc$ for some $c\neq a$ (otherwise $xa$ would be internal). Because $xa$ is a supermaximal right-extension, the set of left-extensions $L_{xa}$ and $L_{yb}\subseteq L_{xc}$ are disjoint by Observation 5.1. Therefore, the selected super-maximal extensions are of the form $\lambda(xa)=ux^{R}\cdot l_{xa}$ and $\lambda(yb)=vy^{R}\cdot l_{xc}$ with $l_{xa}\neq l_{xc}$, thus, distinct.
@@
    Similarly, if $x=y$, the two super-maximal extensions $xa$ and $xb$ have disjoint sets of left-extensions $L_{xa}$ and $L_{xb}$, and hence $\lambda(xa)\neq\lambda(xb)$ because they end in different characters.
@@
    Let $yb$ be the child of $xa$. Let $c\neq b$ be such that xazcxazc is a right-extension in $w$, and $\lambda(xa)=uz^{R}ax^{R}\cdot l_{yc}$ with $l_{yc}\in L_{yc}$. Note that $\lambda(xa)$ differs from $\lambda(yb)$ because $\lambda(yb)$ finishes with some $l_{yb}\in L_{yb}$ and, by Observation 5.1, $L_{yb}\cap L_{yc}=\emptyset$.
@@
    If $yb$ is a farther descendant of $xa$, then xazxaz has children by $c$ and by some $d\neq c$ where xazdxazd prefixes $y$. The string xazdxazd is a supermaximal right-extension because there is a node for it in the tree, hence, by Observation 5.1, it holds $L_{xazc}\cap L_{xazd}=\emptyset$. As $L_{yb}\subseteq L_{xazd}$, it follows $L_{xazc}\cap L_{yb}=\emptyset$. Therefore, $\lambda(xa)$ and $\lambda(yb)$ end in different symbols.
@@
Figure 3: Case 4) and 5) of Lemma 10, where $xa$ is an internal node of the PSR tree, ybyb is the only child of $xa$, and $y^{\prime}d$ is a farther descendant of $xa$ (if it exists). Note xazcxazc is a right-extension, but not super-maximal (otherwise, $xa$ would not be unary). In this case, $\lambda(xa)$ is suffixed by $y^{R}\cdot l_{yc}$, whereas $\lambda(yb)$ and $\lambda(y^{\prime}d)$ are suffixed by $y^{R}\cdot l_{yb}$ and $y^{R}\cdot l_{y^{\prime}d}$, respectively. Note how $yb$ and all its descendants yield $\lambda(\cdot)$ values differing from $\lambda(xa)$ in their last symbol.

### v5

@@ 4.2 Sensitivity to edit operations and rotations
    $\texttt{b}^{4m+2}\texttt{a}$ and $\texttt{b}^{4m+3}$,
  5. 5.
    $\texttt{b}^{m-1}\texttt{a}\texttt{b}^{m-1}\texttt{a}$ and $\texttt{b}^{m-1}\texttt{a}\texttt{b}^{m}$.
Thus, we have $\mathtt{sre}(w_{m}^{\prime})=4m-2$.
Deletions. Let us consider the string $w_{m}^{\prime}$ obtained by adding a just following the second a of $w_{m}$. Same as above, $\texttt{b}^{2m-k+1}\texttt{a}\texttt{b}^{k}$ occurs only once for each $k\in[1..m-1]$, and any other occurrences of $\texttt{b}^{i}\texttt{a}\texttt{b}^{j}$ disappears when $i+j>2m$. This is almost as the previous case, but we have to consider the right-extension of $\texttt{b}^{2m}\texttt{a}$ found twice, once in the prefix $v^{\prime}=\texttt{a}\texttt{b}^{2m}\texttt{a}\texttt{a}\texttt{b}^{2m+2}$ and once across $v^{\prime}u_{1}$, whereas $\texttt{a}\texttt{b}^{2m}\texttt{a}$ and $\texttt{a}\texttt{b}^{2m+1}$ still remain supermaximal right-extensions. We then have that the elements of $\mathcal{S}_{r}(w_{m}^{\prime})$ are the following:
@@
    $\texttt{a}\texttt{b}^{2m}\texttt{a}$ and $\texttt{a}\texttt{b}^{2m+1}$,
@@
    $\texttt{a}\texttt{b}^{2m-1}\texttt{a}$ and $\texttt{a}\texttt{b}^{2m}$,
@@
    $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k-1}\texttt{a}$ and $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}$ (for $k\in[1,m-2]$),
@@
    $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}$ and $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k+1}$ (for $k\in[1,m-2]$),
  5. 5.
    $\texttt{b}^{2m}\texttt{a}\texttt{a}$ and $\texttt{b}^{2m}\texttt{a}\texttt{b}$,
  6. 6.
    $\texttt{b}^{2m+1}\texttt{a}$ and $\texttt{b}^{2m+2}$,
  7. 7.
    $\texttt{b}^{m-1}\texttt{a}\texttt{b}^{m-1}\texttt{a}$ and $\texttt{b}^{m-1}\texttt{a}\texttt{b}^{m}$.
Thus, we have $\mathtt{sre}(w_{m}^{\prime})=4m+2$.
Rotations. Lastly, we consider the string $w_{m}^{\prime}$ that is a rotation of $w_{m}$ beginning in the second a. Note that $u_{m}$ is now followed by $\texttt{a}\texttt{b}^{2m}$. Hence, with respect to $w_{m}$, the substring $\texttt{b}^{m+1}\texttt{a}\texttt{b}^{m-1}\texttt{a}\texttt{b}^{m+1}\texttt{a}\texttt{b}^{m-1}$ occurs twice in $w^{\prime}_{m}$. Thus, we have that the elements of $\mathcal{S}_{r}(w_{m}^{\prime})$ are the following:
@@
    $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k-1}\texttt{a}$ and $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}$ (for $k\in[1,m-1]$),
@@
    $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}$ and $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k+1}$ (for $k\in[1,m-1]$).
Thus, we have $\mathtt{sre}(w_{m}^{\prime})=4m-2$. ∎
Corollary 2.
@@
There exists a string family where $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})\in\Omega(\sqrt{n})$.
Proof.
The strings $w_{m}$ in Lemmas 6 and 7 have length $\Theta(m^{2})$. Since $\mathtt{sre}(w_{m})=6m+4$ and $\mathtt{sre}(w_{m}^{\prime})\leq 4m+2$ for the string $w_{m}^{\prime}$ that results from applying any edit operation or rotation on $w_{m}$, the result follows. ∎
4.3 Sensitivity to string reversal
We now show that $\mathtt{sre}$ can grow by $\Omega(n)$ upon string reversals. We remind that the maximum additive growth is always $O(n)$ because $\chi=O(n)$.
Lemma 8.
There exists a string family where $\mathtt{sre}(w^{R})-\mathtt{sre}(w)=n/5-1$.
Proof.
Such a family is composed of the strings $w_{k}=\prod_{i=1}^{k}\texttt{c}\texttt{a}_{i}\text{\tt\#}_{i}\texttt{a}_{i}\text{\tt\textdollar}_{i}$ on the alphabet $\Sigma=\bigcup_{i=1}^{k}\{\texttt{a}_{i},\text{\tt\#}_{i},\text{\tt\textdollar}_{i}\}\cup\{\texttt{c}\}$. The size of $w_{k}$ is $n=5k$.
Observe that any substring of $w_{k}$ containing $\text{\tt\#}_{i}$ or $\text{\tt\textdollar}_{i}$ is not right-maximal because those symbols occur once. Moreover, every substring of length $2$ of $w_{k}$ is also unique and not right-maximal. Therefore, all the supermaximal extensions in $w_{k}$ and $w_{k}^{R}$ have length at most $2$.
One can verify that the set of supermaximal extensions of $w_{k}$ is $\mathcal{S}_{r}(w_{k})=\bigcup_{i=1}^{k}\{\texttt{a}_{i}\text{\tt\#}_{i},\texttt{a}_{i}\text{\tt\textdollar}_{i},\texttt{c}\texttt{a}_{i}\}\cup\{\texttt{c}\}$. Thus, $\mathtt{sre}(w_{k})=3k+1$.
On the reversed string, one can verify that the set of supermaximal extensions of $w_{k}^{R}$ is $\mathcal{S}_{r}(w_{k}^{R})=\bigcup_{i=1}^{k}\{\texttt{a}_{i}\texttt{c},\texttt{a}_{i}\text{\tt\#}_{i},\texttt{a}_{i}\}\cup\bigcup_{i=1}^{k-1}\{\texttt{c}\text{\tt\textdollar}_{i}\}\cup\{\text{\tt\textdollar}_{k}\}$. Thus, $\mathtt{sre}(w_{k}^{R})=4k$, and the claim follows. ∎
5 Multiplicative Sensitivity of $\chi$ to String Operations
We now focus our attention to multiplicative sensitivity. We first show that $\chi$ has a multiplicative sensitive to reversals that is a constant strictly larger than $1$. Then we show a relation between the multiplicative sensitivities of the other operations. We finally express both additive and multiplicative sensitivities in terms of the substring complexity $\delta$.
5.1 Sensitivity to edits and rotations
A simple consequence of previous results is that the multiplicative sensitivity of edits and rotations is at least 3/2o(1)3/2-o(1).
Corollary 3.
There exists a string family where $\lim_{n\rightarrow\infty}\mathtt{sre}(w)/\mathtt{sre}(w^{\prime})=3/2$ upon edit operations or rotations.
Proof.
The strings $w_{m}$ in Lemmas 6 and 7 satisfy $\mathtt{sre}(w_{m})=6m+4$, whereas $\mathtt{sre}(w_{m}^{\prime})\leq 4m+2$ holds for the string $w_{m}^{\prime}$ that results from applying any edit operation or rotation on $w_{m}$. The result follows. ∎
We conjecture that the multiplicative sensitivity of all those operations is indeed constant. The following results relate them all in this sense.
Lemma 9.
For any $\mathtt{op}_{1},\mathtt{op}_{2}\in\{\mathtt{ins},\mathtt{del},\mathtt{sub},\mathcal{R}\}$, $\mathtt{sre}(\mathtt{op}_{1}(w))/\mathtt{sre}(w)=\Theta(1)$ if and only if $\mathtt{sre}(\mathtt{op}_{2}(w))/\mathtt{sre}(w)=\Theta(1)$.
Proof.
Suppose $\mathtt{sre}(w^{\prime})/\mathtt{sre}(w)=\Theta(1)$ for any $w=xy$ and $w^{\prime}=yx$ a rotation of $w$. Then, writing $X\approx Y$ for $X=\Theta(Y)$, it holds
$\mathtt{sre}(xcy)\approx\mathtt{sre}(yxc)\approx\mathtt{sre}(yx)\approx\mathtt{sre}(xy)$
and
$\mathtt{sre}(xcy)\approx\mathtt{sre}(yxc)\approx\mathtt{sre}(yxd)\approx\mathtt{sre}(xdy),$
which implies $\mathtt{sre}(w^{\prime})/\mathtt{sre}(w)=\Theta(1)$ for all edit operations. This holds because appending symbols change $\mathtt{sre}$ by $O(1)$ only.
Now assume $\mathtt{sre}(w^{\prime})/\mathtt{sre}(w)=\Theta(1)$ for any $w=xy$ and $w^{\prime}=xcy$ (insertion) or vice versa (deletion). Then, using a \$ as in Lemma 5,
$\mathtt{sre}(xy)\approx\mathtt{sre}(x\text{\tt\textdollar}y)\leq\mathtt{sre}(yx\text{\tt\textdollar})=O(\mathtt{sre}(yx)),$
$\mathtt{sre}(yx)\approx\mathtt{sre}(y\text{\tt\textdollar}x)\leq\mathtt{sre}(xy\text{\tt\textdollar})=O(\mathtt{sre}(xy)),$
and thus $\mathtt{sre}(xy)\approx\mathtt{sre}(yx)$. Finally, assume $\mathtt{sre}(w^{\prime})/\mathtt{sre}(w)=\Theta(1)$ for any $w=xcy$ and $w^{\prime}=xdy$ (substitution). Then
$\mathtt{sre}(xcyd)\approx\mathtt{sre}(x\text{\tt\textdollar}yd)\leq\mathtt{sre}(ydx\text{\tt\textdollar})=O(\mathtt{sre}(ydxc)),$
$\mathtt{sre}(ydxc)\approx\mathtt{sre}(y\text{\tt\textdollar}xc)\leq\mathtt{sre}(xcy\text{\tt\textdollar})=O(\mathtt{sre}(xcyd)),$
thus $\mathtt{sre}(xcyd)\approx\mathtt{sre}(ydxc)$, which includes every rotation. ∎
5.2 Sensitivity to string reversals
The proof of Lemma 8 immediately yields a lower bound of 4/3o(1)4/3-o(1).
Corollary 4.
There exists a string family where $\lim_{n\rightarrow\infty}\mathtt{sre}(w^{R})/\mathtt{sre}(w)=4/3$.
Proof.
For the family described in the proof of Lemma 8, it holds that $\mathtt{sre}(w_{k})=3k+1$ and $\mathtt{sre}(w_{k}^{R})=4k$. ∎
[…]
@@ 5.2 Sensitivity to string reversals
In the following, we denote the set of symbols preceding a substring $x$ (or left-extensions of $x$) of $w$ by $L_{x}(w)$, and we use $l_{x}(w)$ to denote an arbitrary element of $L_{x}(w)$ (we drop $w$ if it is clear from the context). Please note that left-extensions are not analogous to right-extensions. A right-extension is a string $xa$ such that $xb$ also occurs for some $b\neq a$, while a left-extension of $x$ is just any symbol preceding $x$. A left-extension can be unique, whereas right-extensions come (at least) in pairs. We will rely on the following observation.
Observation 1.
If $xa$ is a supermaximal right-extension, then $L_{xa}\cap L_{xb}=\emptyset$ for any $b\neq a$.
Proof.
@@
Otherwise, $xa$ is not supermaximal, because there exists $c\in L_{xa}\cap L_{xb}$, for which $cxa\in\mathcal{F}_{w}$ and $cxb\in\mathcal{F}_{w}$, and thus $xa$ is a proper suffix of the right-extension cxacxa of the right-maximal string $cx$. ∎
@@
Let us introduce a special tree of supermaximal extensions, which will be helpful to prove our result.
Definition 6.
@@
The prefix supermaximal right-extension tree (PSR tree) of $w$ is a compacted trie of all the supermaximal right-extensions of $w$. Since some of those strings can be prefixes of others, internal nodes of the trie may also represent strings, and their corresponding nodes are not compacted. Distinct leaves do not represent prefixes of each other.
The following result is key to our proof.
Lemma 10.
@@
Let $w$ be a string starting with a unique symbol \$. If $x$ is a right-maximal substring of $w$, and $xa$ is a supermaximal right-extension of $x$ in $w$, then $x^{R}$ is a right-maximal substring of $w^{R}$.
Proof.
@@
Because $xa$ is a right-extension in $w$, there exists $xb\in\mathcal{F}_{w}$ with $b\neq a$. Any symbol preceding $xa$ cannot precede $xb$, otherwise, $xa$ would not be supermaximal. Since $w$ starts with the unique symbol \$, if $x\neq\varepsilon$, then $xa$ and $xb$ have occurrences that are not prefixes of $w$, and hence, they are preceded by different symbols. It follows $x^{R}$ is right-maximal in $w^{R}$. ∎
Lemma 11.
@@
    If $x$ is a proper prefix of $y$, then and $y$ has to start with $xc$ for some $c\neq a$ (otherwise $xa$ would be internal). Because $xa$ is a supermaximal right-extension, the set of left-extensions $L_{xa}$ and $L_{yb}\subseteq L_{xc}$ are disjoint by Observation 1. Therefore, the selected supermaximal extensions are of the form $\lambda(xa)=ux^{R}\cdot l_{xa}$ and $\lambda(yb)=vy^{R}\cdot l_{xc}$ with $l_{xa}\neq l_{xc}$, thus, distinct.
@@
    Similarly, if $x=y$, the two supermaximal extensions $xa$ and $xb$ have disjoint sets of left-extensions $L_{xa}$ and $L_{xb}$, and hence $\lambda(xa)\neq\lambda(xb)$ because they end in different characters.
@@
    Let $yb$ be the child of $xa$. Let $c\neq b$ be such that xazcxazc is a right-extension in $w$, and $\lambda(xa)=uz^{R}ax^{R}\cdot l_{yc}$ with $l_{yc}\in L_{yc}$. Note that $\lambda(xa)$ differs from $\lambda(yb)$ because $\lambda(yb)$ finishes with some $l_{yb}\in L_{yb}$ and, by Observation 1, $L_{yb}\cap L_{yc}=\emptyset$.
@@
    If $yb$ is a farther descendant of $xa$, then xazxaz has children by $c$ and by some $d\neq c$ where xazdxazd prefixes $y$. The string xazdxazd is a supermaximal right-extension because there is a node for it in the tree, hence, by Observation 1, it holds $L_{xazc}\cap L_{xazd}=\emptyset$. As $L_{yb}\subseteq L_{xazd}$, it follows $L_{xazc}\cap L_{yb}=\emptyset$. Therefore, $\lambda(xa)$ and $\lambda(yb)$ end in different symbols.
@@
Figure 4: Case 4) and 5) of Lemma 11, where $xa$ is an internal node of the PSR tree, ybyb is the only child of $xa$, and $y^{\prime}d$ is a farther descendant of $xa$ (if it exists). Note xazcxazc is a right-extension, but not supermaximal (otherwise, $xa$ would not be unary). In this case, $\lambda(xa)$ is suffixed by $y^{R}\cdot l_{yc}$, whereas $\lambda(yb)$ and $\lambda(y^{\prime}d)$ are suffixed by $y^{R}\cdot l_{yb}$ and $y^{R}\cdot l_{y^{\prime}d}$, respectively. Note how $yb$ and all its descendants yield $\lambda(\cdot)$ values differing from $\lambda(xa)$ in their last symbol.
5.3 Sensitivity as a function of $\delta$
Lemma 12.
Proof.